# Networked Multiagent Reinforcement Learning for Peer-to-Peer Energy Trading

### 25 Jan 2024 v1

@@ Introduction
Objective function: At each time step $t$, agent $i$ receives the system state $s_{t}$, chooese an action $a_{i,t}\in\mathcal{A}_{i}$, and receives a reward $r_{i,t}$. At time $t$, agent $i$’s decision, given the system state $s_{t}\in\mathcal{S}=\prod_{i=1}^{I}\mathcal{S}_{i}$, is determined by the policy $\pi_{\theta_{i}}(\cdot|s)$. This policy is a density function mapping from $\mathcal{A}_{i}$ to [0, $\infty$), and it is parameterized by $\theta_{i}$. Here, $\theta_{i}$ is an element of a generic set $\Theta_{i}\in\mathcal{R}^{m_{i}}$, where the dimensionality $m_{i}$ is determined by each agent. In choosing $m_{i}$, each agent faces a tradeoff: it should be large enough to avoid underfitting, yet not so large as to risk overfitting and increased training time. Let $\theta=[\theta_{1},\cdots,\theta_{I}]$ be the concatenation of $\theta_{i}$, and $\pi_{\theta}(\cdot|s)=\prod_{i=1}^{I}\pi_{\theta_{i}}(\cdot|s)$ be the joint policy of all agents. In the following discussion of this section, we assume that the Markov chains $\left\{s_{t}\right\}_{t\geq 0}$ induced by every agents’ collective policy $\pi_{\theta}$ where $\theta\in\Theta$ is ergodic with a stationary distribution $\rho_{\theta}$.
@@
to optimize the average system long-term reward. The objective function resembles a cooperative game, which may face similar critiques regarding the need for each agent to access system-wide state variables in a decentralized setting. A pertinent response, as previously mentioned, is the use of control automation: the objective function can be integrated into intelligent control devices, making them ‘black boxes’ for end-users. Additionally, in developing a MARL framework with provable convergence, some level of global information sharing among agents seems necessary. The collective objective function, as outlined in (9), is essential for decentralized agents to reach a consensus.
@@
For each agent $i=1,\ldots,I$, since the market clearing price $P_{t}$ and the agent’s bid $b_{i,t}$ are bounded above (by their PV and battery capacities), according to Eq. (III) and (3), respectively, the first term in an agent’s reward function, $R^{m}_{i}$, is then bounded above based on Eq. (6), for any $a\in\mathcal{A}:=\Pi_{i=1}^{I}\mathcal{A}_{i}$ and $s\in\mathcal{S}$. The second term, $R^{v}$, is bounded above by 0 based on its definition in Eq. (8). Hence, agent $i$’s reward $r_{i,t}$ is bounded at any $t$, and for any $a\in\mathcal{A}$ and $s\in\mathcal{S}$. Since the number of agents is finite, there exists a common upper bound for all $r_{i,t}$’s, and let us denote it as $\overline{R}$. Then by the formulation in (9), $J(\theta)\leq\overline{R}$ for all $\theta\in\Theta$. By the well-known least-upper-bound property of real numbers, we know that $\sup_{\theta}J(\theta)$ exists and is finite. ∎
@@
The function $Q_{\theta}(s,a)$ cannot be calculated by each agent, even if the global state and action information are shared, since the joint policy distribution $\pi_{\theta}$ is not known. Instead, each agent uses $\hat{Q}(\cdot,\cdot;\omega^{i}):\mathcal{S}\times\mathcal{A}\to\mathcal{R}$, which is a class of function parametrized by $\omega^{i}\in\mathcal{R}^{K}$, where $K$ is the dimension of the $\omega^{i}$, to approximate the action-value function $Q_{\theta}(s,a)$. Note that unlike the parameters $\theta_{i}$, where agents can choose their own policy models (and hence different dimensions of $\theta_{i}$), the dimension of $\omega^{i}$ has to be the same across all agents (and hence $K$ does not have an agent index) to facilitate the consensus update.
@@
where $\Gamma^{i}:\mathbb{R}^{m_{i}}\rightarrow\Theta_{i}\subset\mathbb{R}^{m_{i}}$ is a projector operator. An example is the orthogonal projector mapping: $\Gamma^{i}(x)\triangleq\underset{y\in\Theta_{i}}{\arg\min}\|y-x\|\,\forall\mathbf{x}\in\mathbb{R}^{m_{i}}$ where $\Theta$ is a compact convex set. The $\widehat{I}_{t}^{Q,i}$ term in (14) is given as follows:
@@
To limit communication among agents, each agent $i$ only shares their local parameter $\widetilde{\omega}_{t}^{i}$ with the others; no other information needs to be shared. A consensual estimate of $Q_{\theta}(s,a)$ can then be estimated through updating the parameter $\omega_{t+1}^{i}$ as follows:
@@
To apply such an algorithm to the P2P energy trading, we illustrate the conceptual framework in Fig. 2. The detailed algorithm is presented in Algorithm 1.
In the next section, we present the conditions under which Algorithm 1 can be shown to converge to a steady state.
@@
The power flow equation of the distribution network is solved, the P2P energy market is cleared, and the information of joint actions $a_{0}=\left(a_{i,0},\ldots,a_{I,0}\right)$ is sent to each agent.
@@
      The power flow equation of the distribution network is solved, the P2P energy market is cleared, and the information of joint actions $a_{t+1}=\left(a_{i,t+1},\ldots,a_{I,t+1}\right)$ is sent to each agent.
@@
In this section, we discuss the convergence results of Algorithm 1. For ease of notation, we drop the $t$ index of the parameters $\theta_{i}$ for $i\in\mathcal{I}$ in the following discussions.
@@
$||P^{n}(x,\cdot)-\pi(\cdot)||\leq Mc^{n},\ n=1,2,3,\ldots$
@@
For any $i\in\{1,\dots,I\},s\in\mathcal{S}$ and $\theta_{i}\in\Theta_{i}$, an agent $i$’s policy, aka the conditional probability density over the action space $\mathcal{A}_{i}$, is positive; that is, $\pi_{\theta_{i}}\left(a_{i}\mid s\right)>0$ for all $a_{i}\in\mathcal{A}_{i}$. Additionally, $\pi_{\theta_{i}}\left(\cdot\mid s\right)$ is assumed to be continuously differentiable with respect to the parameter $\theta_{i}$ over $\Theta_{i}$. Finally, the Markov chains $\left\{s_{t}\right\}_{t\geq 0}$ and $\left\{\left(s_{t},a_{t}\right)\right\}_{t\geq 0}$ induced by the agents’ collective policies $\pi_{\theta}$ are both geometrically ergodic, with the coressponding stationary distribution denoted by $\rho_{\theta}$ and $\tilde{\rho_{\theta}}$, respectively.
Remark 1 The first part of Assumption 1 is a standard assumption on the policy function. One example of such a policy is Gaussian, i.e., $\pi_{\theta_{i}}(\cdot\mid s)\sim$ $\mathcal{N}\left(\eta_{\theta_{i}}(s),\Sigma_{i}\right)$, where $\eta_{\theta_{i}}(s):\mathcal{S}\rightarrow\mathcal{A}_{i}\in\mathbb{R}^{n}$ is a fully connected neural network parametrized by $\theta_{i}$ and is continuously differentiable with respect to $\theta_{i}$ and $s$, and $\Sigma_{i}\in\mathbb{R}^{n\times n}$ is the covariance matrix.
@@
where $[s,a]$ is the concatenation of $s$ and $a$, $\gamma_{j}\in\mathcal{R}^{+}$ for $j=1,\cdots,K$, $c_{j}\in\mathcal{R}^{|\mathcal{S}|+|\mathcal{A}|}$ for $j=1,\cdots,K$. $\gamma_{j}$ and $c_{j}$ can be chosen arbitrarily, as long as $(\gamma_{j},c_{j})$ are different with different $j$ so that the feature functions are linearly independent.
@@
Assume that each agent’s state space $\mathcal{S}_{i}$ is compact for $i=1,\ldots,I$. Under Assumptions 1 – 4, the sequences $\left\{\theta_{i,t}\right\}$, $\left\{\mu_{t}^{i}\right\}$ and $\left\{\omega_{t}^{i}\right\}$ generated from Algorithm 1 satisfy the following.
@@
The general convergence result of the consensus algorithm is proved in [1] under six assumptions. Our Assumption 1, 2, 3 and 4 directly correspond to four assumptions in [1]. The fifth assumption [1] regarding agents’ time-varying neighborhood is trivially true in our setting since we assume agents can communicate through the whole network and the network topology does not change over time. Hence, the only remaining thing to show is the uniform boundedness of agents’ reward $r_{i,t}$.
By the definition of the market clearing price $P_{t}$ in (III), it is bounded between $FIT$ and $UR$. Together with the compactness assumption of the state space $\mathcal{S}_{i}$ (which is reasonable since an agent’s energy demand and battery/PV capacities are all bounded), the $R_{i}^{m}$ component in the reward function, as defined in (5), is uniformly bounded. The other term, the voltage violation penalty $R_{i}^{v}$, is uniformly upper bounded by 0 and lower bounded by $-\lambda NM$, based on its definition. Hence, $r_{i,t}$ is uniformly bounded for all $i$ and $t$, and the convergence results follow directly from Theorem 2 and 3 in [1]. ∎
Remark 4. While the theorem establishes the convergence to a steady state of the critic and actor steps, it does not address how good the system reward is upon convergence. A key concern in this context is the potential inadequacy of the linear approximation of the critic function $\hat{Q}(s,a;\omega)$ in Assumption 3. In our model, the distribution network encompasses a multitude of agents, each exhibiting non-linear behaviors and interactions that may not be adequately captured by simple linear models. One natural idea is to deep neural networks (DNNs) instead, as they excel in capturing and modeling non-linear relationships. Specifically, the layered architecture of DNNs allows for the abstraction and hierarchical representation of data, enabling the network to learn rich and intricate representations of the environment. This is particularly crucial in our context, where the environment is characterized by a high degree of stochasticity and dynamical complexity. The downside of using DNNs, however, is that the proof of Theorem IV.1 is no longer applicable. Consequently, the convergence of the critic and actor steps remains an unresolved issue, meriting further investigation in future research.
@@
We test our problem using the IEEE 13-bus test feeder. In our simulations, we set each bus to have one prosumer, except for the substation bus; hence, there are 12 prosumers (aka agents). The length of each time step is set to be 1 hour. The distribution network’s voltage limit is set to $[0.96\ pu,1.04\ pu]$. The voltage violation penalty, $M$, is set at $10^{4}$. The price ceiling and floor, that is, the utility rate $UR$ and feed-in tariff $FIT$, are set at 14 ¢/kWh and 5 ¢/kWh, respectively. The capacity of each agent’s PV is set to be 30kW, and the energy storage capacity is set to be 50kWh, with the charging and discharging efficiency being 0.95 and 0.9 for each agent. The PV generation of the prosumers and the total base load of the pure consumers on each bus have fixed diurnal shapes. These shapes represent the mean values at the corresponding hour of the day. The detail of these shapes can be found in [2] due to the page limit. The power flow equations (7) are solved by the open-source distribution system simulator OpenDSS [25].
@@
where $\sigma^{(1)}_{\theta_{i}}(s)$ represents the mean value of the charging/discharging action, $\sigma^{(2)}_{\theta_{i}}(s)$ denotes the mean value of the smart inverter action. Similarly, $\sigma^{(3)}_{\theta_{i}}(s)$ and $\sigma^{(4)}_{\theta_{i}}(s)$ correspond to the variances of the charging/discharging action and the smart inverter action, respectively. The function $\sigma_{\theta_{i}}(s)=\begin{bmatrix}\sigma^{(1)}_{\theta_{i}}(s)\\ \cdots\\ \sigma^{(4)}_{\theta_{i}}(s)\end{bmatrix}:\mathcal{S}\to\mathcal{R}^{4}$ is implemented as a DNN, parameterized by $\theta_{i}$. This network comprises two fully connected layers, each consisting of 256 neurons, and employs a $tanh$ activate function following each layer. Specifically,
$\sigma_{\theta_{i}}(s)=$
${\theta^{(3)}_{i}}tanh\Big{(}{\theta^{(2)}_{i}}tanh({\theta^{(1)}_{i}}s+{\theta^{(1)}_{i}}^{bias}+{\theta^{(2)}_{i}}^{bias}\Big{)}+{\theta^{(3)}_{i}}^{bias},$ (27)
where $\theta_{i}=({\theta^{(1)}_{i}},{\theta^{(1)}_{i}}^{bias},{\theta^{(2)}_{i}},{\theta^{(2)}_{i}}^{bias},{\theta^{(3)}_{i}},{\theta^{(3)}_{i}}^{bias})$ are the policy parameters, ${\theta^{(1)}_{i}}$ is a $256\times|\mathcal{S}|$ matrix, ${\theta^{(2)}_{i}}$ is a $256\times 256$ matrix, ${\theta^{(3)}_{i}}$ is a $4\times 256$ matrix, ${\theta^{(1)}_{i}}^{bias}$ is a $|\mathcal{S}|\times 1$ vector, ${\theta^{(2)}_{i}}^{bias}$ is a $256\times 1$ vector, ${\theta^{(3)}_{i}}^{bias}$ is a $4\times 1$ vector, and for any vector $x\in\Re^{n}$,
$tanh(x):=\begin{bmatrix}\frac{e^{x_{1}}-e^{-x_{1}}}{e^{x_{1}}+e^{-x_{1}}}\\ \cdots\\ \frac{e^{x_{1}}-e^{-x_{1}}}{e^{x_{n}}+e^{-x_{n}}}\end{bmatrix},$ (31)
where $x_{j}$, $j=1,\ldots,n$, is the $j$-th element of $x$.
A similar neural network architecture is also applied to the approximator of the action-value function $\hat{Q}(\cdot,\cdot;\omega^{i})$. The settings of stepsizes are $\beta_{\omega,t}=1/t^{0.65}$ and $\beta_{\theta,t}=1/t^{0.85}$.
@@
We compare the consensus MARL with two other algorithm frameworks: (i) a fully decentralized framework in which each agent just solves their own reinforcement learning problem using the PPO algorithm [2], while completely ignoring multiagent interactions, and (ii) the MADDPG approach, originally proposed in [4].
@@
where $\beta$ is the step size and $\widehat{\nabla J(\theta_{t})}$ is a stochastic estimate approximating the gradient of the performance measure. The three MARL approaches differ in how each agent updates $\widehat{\nabla J(\theta_{t})}$. To highlight such differences, we present the specific gradient approximations for each of the three approaches below.
In the fully decentralized approach, the gradient update uses local states and actions only. Particularly, for each $i=1,\ldots,I$,
@@
In the MADDPG approach, each agent makes decisions based on only their own state variables. However, their action-value functions are updated based on the global states and actions of all agents by a central authority, who is assumed to have access to the global information. Specifically, for each agent $i=1,\ldots,I$, the gradient approximation in MADDPG is
@@
While the action-value function $Q_{i}^{\pi}\left(\bm{s};\bm{a_{1},\ldots,a_{I}}\right)$ is the same as in (35), the key difference lies in the policy function. As stated in Section IV-A and reflected in (37), each agent $i$’s policy depends on all agents’ states $\bm{s}$. The other key difference between the consensus MARL and the MADDPG algorithm is how the action-value function $Q_{i}^{\pi}$ is updated. In the MADDPG framework, the function is updated through centralized training through a central authority; while in the consensus MARL, updates are executed via a decentralized consensus process, as described in Section IV-B.
@@
Figure 3 displays the convergence curves of the 30-episode moving average of the episodic total cost in the distribution network for the three different MARL algorithms.
@@
It is evident that the mean episodic total reward gradually converges to a high level at the end of the training in both frameworks, indicating the effective convergence of the consensus-update framework. Furthermore, the agents in the consensus-update framework achieve a higher stationary reward level than those in the fully decentralized framework, which is expected since the consensus-update framework benefits from improved approximation of the action-value function through communication. The reward under MADDPG is higher than that under the fully decentralized PPO, but lower than that under consensus-update framework. On one hand, MADDPG has the process of centralized learning which can guarantee a better performance than the fully decentralized one. However, there is no communication among agents in the process of decentralized execution in MADDPG, so each agent’s decision-making is only based on their own local states, which leads to a worse overall performance than the consensus-update framework.
Figure 4 presents the market price of the P2P market in both frameworks in chronological order over the last three days after thousands of training episodes.
@@
Notably, the market price in the consensus-update framework demonstrates a tendency to be higher during off-peak hours and lower during peak hours when compared to the fully decentralized framework. This indicates that agents employing the consensus-update algorithm have developed more sophisticated trading strategies compared to those using the fully decentralized PPO algorithm. Specifically, the consensus-algorithm-induced strategies involve buying more and selling less during off-peak hours, and conversely, selling more and buying less during peak hours, when compared with the strategies produced by the purely decentralized algorithm.
Figure 5 shows the gradual reduction of the system voltage deviation after training.
@@
It can be seen that the voltage violations of the system are all high for the three frameworks at the beginning. However, that of the consensus framework drops and converges to zero quickly at the fastest rate, and the other two converge to almost zero as well after half of the training process. On the one hand, the results are remarkable in the sense that voltage violation reduction is realized through purely decentralized learning for all three frameworks, while the consensus approach has the best performance.
@@
This study has developed a market and algorithmic framework to enable energy consumers’ and prosumers’ participation in local P2P energy trading. Utilizing reinforcement learning algorithms, we have automated bidding for agents while ensuring decentralized decision-making. The SDR-based market clearing addresses challenges with zero-marginal-cost resources and simplifies bidding. Additionally, our MARL framework includes voltage constraints of the physical network, setting a foundation towards real-world applications.
However, this research represents just the initial phase in the practical application of a P2P energy trading market, with several challenges ahead. Theoretically, the scalability of the MARL framework is crucial; if the scalability of the consensus MARL algorithm is limited, a mean-field approach, as in [28], could be considered, where agents operate based on a mean-field equilibrium. Practically, while the current model uses discrete, synchronous trading rounds, future work should investigate continuous and asynchronous trading among agents.

### 27 Jan 2024 v2

@@ Introduction
Objective function: At each $t$, agent $i$ receives the system state $s_{t}$, chooeses an action $a_{i,t}\in\mathcal{A}_{i}$, and receives a reward $r_{i,t}$. Agent $i$’s decision, given the system state $s_{t}\in\mathcal{S}$, is determined by the policy $\pi_{\theta_{i}}(\cdot|s_{t})$. This policy is a density function mapping from $\mathcal{A}_{i}$ to [0, $\infty$) and is parameterized by $\theta_{i}$. Here, $\theta_{i}$ is an element of a generic set $\Theta_{i}\in\mathcal{R}^{m_{i}}$, where the dimensionality $m_{i}$ is determined by each agent. In choosing $m_{i}$, each agent faces a tradeoff: it should be large enough to avoid underfitting, yet not too large to risk overfitting and increased training time. Let $\theta:=[\theta_{1},\cdots,\theta_{I}]$ and $\pi_{\theta}(\cdot|s)=\prod_{i=1}^{I}\pi_{\theta_{i}}(\cdot|s)$ be the joint policy of all agents. In the following discussion of this section, we assume that the Markov chains $\left\{s_{t}\right\}_{t\geq 0}$ induced by the collective policy $\pi_{\theta}$ is ergodic with a stationary distribution $\rho_{\theta}$.
@@
which is to optimize the average system long-term reward. The objective function resembles a cooperative game, which may face similar critiques regarding the need for each agent to access system-wide state variables in a decentralized setting. A pertinent response, as previously mentioned, is the use of control automation: the objective function can be integrated into intelligent control devices, making them ‘black boxes’ for end-users. Additionally, in developing a MARL framework with provable convergence, some level of global information sharing among agents is likely unaoidable. The collective objective function, as outlined in (9), is essential for decentralized agents to reach a consensus.
@@
Since the market clearing price $P_{t}$, the supply-demand ratio $SDR_{t}$, and agents bid $b_{i,t}$ for $i=1,\ldots,I$ are all bounded above, the first term in an agent’s reward function, $R^{m}_{i}$, is then bounded above based on Eq. (5), for any $a\in\mathcal{A}:=\Pi_{i=1}^{I}\mathcal{A}_{i}$ and $s\in\mathcal{S}$. The second term, $R^{v}$, is bounded above by 0 based on its definition in Eq. (8). Hence, agent $i$’s reward $r_{i,t}$ is bounded at any $t$, and for any $a\in\mathcal{A}$ and $s\in\mathcal{S}$. Since the number of agents is finite, there exists a common upper bound for all $r_{i,t}$’s, and let us denote it as $\overline{R}$. Then by the formulation in (9), $J(\theta)\leq\overline{R}$ for all $\theta\in\Theta$. By the well-known least-upper-bound property of real numbers, we know that $\sup_{\theta}J(\theta)$ exists and is finite. ∎
@@
The function $Q_{\theta}(s,a)$ cannot be calculated by each agent, even if the global state and action information are shared, since the joint policy distribution $\pi_{\theta}$ is not known. Instead, each agent uses $\hat{Q}(\cdot,\cdot;\omega^{i}):\mathcal{S}\times\mathcal{A}\to\mathcal{R}$, which is a class of functions parametrized by $\omega^{i}\in\mathcal{R}^{K}$, where $K$ is the dimension of the $\omega^{i}$, to approximate the action-value function $Q_{\theta}(s,a)$. Note that unlike the parameters $\theta_{i}$, where agents can choose their own policy models (and hence different dimensions of $\theta_{i}$), the dimension of $\omega^{i}$ has to be the same across all agents (and hence $K$ does not have an agent index) to facilitate the consensus update.
@@
where $\Gamma^{i}:\mathbb{R}^{m_{i}}\rightarrow\Theta_{i}\subset\mathbb{R}^{m_{i}}$ is a projector operator. An example is the orthogonal projector mapping: $\Gamma^{i}(x)\triangleq\underset{y\in\Theta_{i}}{\arg\min}\|y-x\|,\ \forall\mathbf{x}\in\mathbb{R}^{m_{i}}$, where $\Theta$ is a compact convex set. The $\widehat{I}_{t}^{Q,i}$ term in (14) is given as follows:
@@
To limit communication among agents, each agent $i$ only shares local parameter $\widetilde{\omega}_{t}^{i}$ with each other; no other information needs to be shared. A consensual estimate of $Q_{\theta}(s,a)$ can then be estimated through updating the parameter $\omega_{t+1}^{i}$ as follows:
@@
To apply such an algorithm to P2P energy trading, we illustrate the conceptual framework in Fig. 2. The detailed algorithm is presented in Algorithm 1.
@@
The power flow equation (7) is solved, the P2P energy market is cleared, and the information of joint actions $a_{0}=\left(a_{i,0},\ldots,a_{I,0}\right)$ is sent to each agent.
@@
      The power flow equation (7) is solved, the P2P energy market is cleared, and the information of joint actions $a_{t+1}=\left(a_{i,t+1},\ldots,a_{I,t+1}\right)$ is sent to each agent.
@@
In this section, we discuss the convergence results of Algorithm 1. For ease of notation, we drop the $t$ index of the parameters $\theta_{i}$ for $i\in\mathcal{I}$ in the following discussions. First, we need a technical condition of Markov chains, known as the geometric ergodicity.
@@
$||P^{m}(x,\cdot)-\pi(\cdot)||\leq Mc^{m},\ m=1,2,3,\ldots$
@@
For each $i\in\mathcal{I},s\in\mathcal{S}$ and $\theta_{i}\in\Theta_{i}$, agent $i$’s policy is positive; that is, $\pi_{\theta_{i}}\left(a_{i}\mid s\right)>0$ for all $a_{i}\in\mathcal{A}_{i}$. Additionally, $\pi_{\theta_{i}}\left(\cdot\mid s\right)$ is assumed to be continuously differentiable with respect to the parameter $\theta_{i}$ over $\Theta_{i}$. Finally, the Markov chains $\left\{s_{t}\right\}_{t\geq 0}$ and $\left\{\left(s_{t},a_{t}\right)\right\}_{t\geq 0}$ induced by the agents’ collective policies $\pi_{\theta}$ are both geometrically ergodic, with the coressponding stationary distribution denoted by $\rho_{\theta}$ and $\tilde{\rho_{\theta}}$, respectively.
Remark 1 The first part of Assumption 1 is a standard assumption on policy functions. One example of such a policy is Gaussian: $\pi_{\theta_{i}}(\cdot\mid s)\sim$ $\mathcal{N}\left(\eta_{\theta_{i}}(s),\Sigma_{i}\right)$, where $\eta_{\theta_{i}}(s):\mathcal{S}\rightarrow\mathcal{A}_{i}\in\mathbb{R}^{n}$ is a fully connected neural network parametrized by $\theta_{i}$ and is continuously differentiable with respect to $\theta_{i}$ and $s$, and $\Sigma_{i}\in\mathbb{R}^{n\times n}$ is the covariance matrix.
@@
where $[s,a]$ is the concatenation of $s$ and $a$, $\gamma_{j}\in\mathcal{R}^{+}$ for $j=1,\cdots,K$, $c_{j}\in\mathcal{R}^{|\mathcal{S}|+|\mathcal{A}|}$ for $j=1,\cdots,K$. The parameters $\gamma_{j}$ and $c_{j}$ can be chosen arbitrarily, as long as $(\gamma_{j},c_{j})$ are different with different $j$ so that the feature functions are linearly independent.
@@
Assume that each agent’s state space $\mathcal{S}_{i}$ is compact for $i\in\mathcal{I}$. Under Assumptions 1 – 4, the sequences $\left\{\theta_{i,t}\right\}$, $\left\{\mu_{t}^{i}\right\}$ and $\left\{\omega_{t}^{i}\right\}$ generated from Algorithm 1 satisfy the following.
@@
The general convergence result of the consensus algorithm is proved in [1] under six assumptions. Our Assumption 1, 2, 3 and 4 directly correspond to four assumptions in [1]. The fifth assumption in [1] regarding agents’ time-varying neighborhood is trivially true here since we assume agents can communicate through the whole network, and the network topology does not change over time. Hence, the only remaining item to show is the uniform boundedness of agents’ reward $r_{i,t}$. By the definition of the market clearing price $P_{t}$ in (III), it is bounded between $FIT$ and $UR$. Together with the compactness assumption of the state space $\mathcal{S}_{i}$ (which is reasonable since an agent’s energy demand and battery/PV capacities are all bounded), the $R_{i}^{m}$ component in the reward function, as defined in (4), is uniformly bounded. The other term, the voltage violation penalty $R_{i}^{v}$, is uniformly upper bounded by 0 and lower bounded by $-\lambda NM$. Hence, $r_{i,t}$ is uniformly bounded for all $i$ and $t$, and the convergence results follow directly from Theorem 2 and 3 in [1]. ∎
Remark 4. While the theorem establishes the convergence to a steady state of the critic and actor steps, it does not address how good the system reward is upon convergence. A key concern in this context is the potential inadequacy of the linear approximation of the critic function $\hat{Q}(s,a;\omega)$ in Assumption 3. In our model, the distribution network encompasses a multitude of agents, each exhibiting non-linear behaviors and interactions that may not be adequately captured by simple linear models. One natural idea is to deep neural networks (DNNs) instead, as they excel in capturing and modeling non-linear relationships. The downside of using DNNs, however, is that the proof of Theorem IV.1 is no longer applicable. Consequently, the convergence of the critic and actor steps remains an unresolved issue, meriting further investigation in future research.
@@
We test the algorithm frameworks using the IEEE 13-bus test feeder [25]. In our simulations, we set each bus to have one prosumer, excluding the substation bus; hence, there are 12 prosumers (and no pure consumers). The length of each time step is set to be 1 hour. The distribution network’s voltage limit is set to $[0.96\ pu,1.04\ pu]$. The amplifying parameter for voltage violation, $\lambda$, is set at $10^{4}$. The price ceiling and floor, that is, the utility rate $UR$ and feed-in tariff $FIT$, are set at 14 ¢/kWh and 5 ¢/kWh, respectively. The capacity of each agent’s PV is set to be 30kW, and the energy storage capacity is set to be 50kWh, with the charging and discharging efficiency being 0.95 and 0.9 for each agent. The PV generation of the prosumers and the total base load of the pure consumers on each bus have fixed diurnal shapes. These shapes represent the average values corresponding to each hour of the day. For a detailed explanation and due to page limitations her, please refer to our earlier paper [2]. Additionally, [2] provides more information on how the random PV output for each agent is generated. The power flow equations (7) are solved by the open-source distribution system simulator OpenDSS [26].
@@
where $\sigma^{(1)}_{\theta_{i}}(s)$ represents the mean value of the charging/discharging action, and $\sigma^{(2)}_{\theta_{i}}(s)$ denotes the mean value of the smart inverter action. Similarly, $\sigma^{(3)}_{\theta_{i}}(s)$ and $\sigma^{(4)}_{\theta_{i}}(s)$ correspond to the variances of the charging/discharging action and the smart inverter action, respectively. The function $\sigma_{\theta_{i}}(s):\mathcal{S}\to\mathcal{R}^{4}$ is implemented as a DNN, parameterized by $\theta_{i}$. This network comprises two fully connected layers, each consisting of 256 neurons, and employs a $tanh$ activate function following each layer. A similar neural network architecture is also applied to the approximator of the action-value function $\hat{Q}(\cdot,\cdot;\omega^{i})$. The settings of stepsizes are $\beta_{\omega,t}=1/t^{0.65}$ and $\beta_{\theta,t}=1/t^{0.85}$.
@@
We compare the consensus MARL with two other algorithm frameworks: (i) a fully decentralized framework in which each agent just solves their own reinforcement learning problem using the PPO algorithm [2], while completely ignoring multiagent interactions, and (ii) the MADDPG approach [4].
@@
where $\beta$ is the step size and $\widehat{\nabla J(\theta_{t})}$ is a stochastic estimate approximating the gradient of the performance measure. The three MARL approaches differ in how each agent updates $\widehat{\nabla J(\theta_{t})}$. To highlight such differences, we present the specific gradient approximations for the three approaches below.
In the fully decentralized approach, the gradient update uses local states and actions only: for each $i\in\mathcal{I}$,
@@
In the MADDPG approach, each agent makes decisions based on only their own state variables. However, their action-value functions are updated based on the global states and actions of all agents by a central authority, who is assumed to have access to the global information. Specifically, for each agent $i\in\mathcal{I}$, the gradient approximation in MADDPG is
@@
While the action-value function $Q_{i}^{\pi}$ is the same as in (30), the key difference lies in the policy function. As stated in Section IV-A and reflected in (32), each agent $i$’s policy depends on all agents’ states $\bm{s}$. The other difference between the consensus MARL and the MADDPG algorithm is how the action-value function $Q_{i}^{\pi}$ is updated. In MADDPG, the function is updated through centralized training by a central authority; while in consensus MARL, updates are executed via a decentralized consensus process, as described in Section IV-B.
@@
Figure 3 displays the convergence curves of the 30-episode moving average of the episodic total reward in the distribution network for the three different MARL algorithms.
@@
It is evident that the mean episodic total reward gradually converges to a high level at the end of the training in all three frameworks. The results show that the consensus MARL algorithm outperforms the fully decentralized framework, which is in line with expectations, since the consensus-update framework benefits from improved approximation of the action-value function through communication. The better performance of consensus MARL over MADDPG, on the other hand, is surprising. Conventionally, one might anticipate a centralized algorithm to outperform decentralized algorithms. The explanation here can be traced to MADDPG’s decentralized execution. In MADDPG, each agent independently updates its actions based on its own local state, despite the centralized nature of the learning process. This approach of relying solely on local information appears to be less effective compared to the consensus update mechanism.
Figure 4 presents the market prices of the P2P market in chronological order over the last three days after thousands of training episodes.
@@
Notably, the market price in the consensus-update framework demonstrates a tendency to be higher during off-peak hours and lower during peak hours. This indicates that agents employing the consensus-update algorithm have developed more sophisticated trading strategies than the other two MARL frameworks. Specifically, the consensus-algorithm-induced strategies involve buying more and selling less during off-peak hours, and conversely, selling more and buying less during peak hours, when compared with the strategies produced by the purely decentralized algorithm.
Figure 5 shows the gradual reduction of the system voltage deviation after training. Initially, it is observed that voltage violations are significantly high across all three frameworks. However, the consensus framework demonstrates a rapid decline in these violations, converging swiftly and effectively to zero at the fastest rate. The remaining two frameworks demonstrate a slower rate of convergence, with the fully decentralized approach seemingly unable to fully eliminate voltage violations.
@@
This study has developed a market and algorithmic framework to enable consumers’ and prosumers’ participation in local P2P energy trading. Utilizing reinforcement learning algorithms, we have automated bidding for agents while ensuring decentralized decision-making. The SDR-based market clearing addresses challenges with zero-marginal-cost resources and simplifies bidding. Additionally, our MARL framework includes voltage constraints of the physical network, setting a foundation for real-world applications.
However, this research represents just the initial phase in the practical application of a P2P energy trading market, with several challenges ahead. Theoretically, the scalability of the MARL framework is crucial; if the scalability of the consensus MARL algorithm is limited, a mean-field approach, as in [28], could be considered, where agents operate with the belief that the market outcome is at a mean-field equilibrium. Practically, while the current model uses discrete, synchronous trading rounds, future work should investigate continuous and asynchronous trading among agents.